when maria was born her mother was 24 years old at present maria's age is 20%of her mother's age how old is maria now?
Answer by mananth(16946)

mother is older than maria by 24 years.
if Maria's age = x
mother's age = x+24
..
x=0.2(x+24)
x=0.2x+4.8
x-0.2x= 4.8
0.8x=4.8
x=4.8/0.8
x=6
Maria's age is 6 years